OpenStudy (anonymous):

@campbell_st I have never done something like this could you be of service?!? Geometric series.

OpenStudy (anonymous):



OpenStudy (anonymous):

@TrojanPoem do you know how to do this?

OpenStudy (anonymous):

add

OpenStudy (anonymous):

put \(i-1\) get \(6^{1-1}=6^0=1\)

OpenStudy (anonymous):

then put a plus sign then put \(i=2\) get \[6^{2-1}=6^1=6\]\

OpenStudy (anonymous):

then put \(i=3\) etc \[1+6+6^2+6^3\]and add them up

OpenStudy (anonymous):

\[\sum_{i=1}^{4}6^{i-1}\] \[=6^{1-1}+6^{2-1}+6^{3-1}+6^{4-1}=6^0+6^1+6^2+6^3=1+6+36+108=?\]

OpenStudy (anonymous):

i would use a calculator

OpenStudy (anonymous):

Thank you @satellite73

OpenStudy (anonymous):

correction 1+6+36+216=?